OpenStudy (anonymous):

square root 3x+7 + square root x+2 = 1

OpenStudy (anonymous):

x^2+3x+2=0

OpenStudy (anonymous):

Can you show the work please. Thank you ^^;

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[\sqrt{3x+7}=1-\sqrt{x+2}\] then suaring on both sides and then solv it

OpenStudy (anonymous):

hve u got it?

OpenStudy (anonymous):

well I squared it and I canceled the squares on the 3x+7 so I am left with square root 3x+7 = (1+square root x+2) squared.

OpenStudy (anonymous):

yah then open rite side

OpenStudy (anonymous):

it is 1-sq root x+2

OpenStudy (mertsj):

sqrt(3x+7)=1-sqrt(x+2) square both sides: 3x+7 = 1 -2sqrt(x=2)+x+2 2x+4 = -2sqrt(x+2) Divide by -2 -x-2=sqrt(x+2) square both sides: x^2+4x+4 = x+2 x^2+3x+2=0 (x+2)(x+1)=0 x=-2, -1 but -1 does not check and is an extraneous root so the answer is -2.

OpenStudy (anonymous):

-2 and -1 both

OpenStudy (mertsj):

-1 does not check since sqrt(-3+7) +sqrt(-1+2) = 2+1=3 NOT 1

OpenStudy (anonymous):

yah ok

OpenStudy (anonymous):

Divide by -2 to both sides?

OpenStudy (mertsj):

of course

OpenStudy (anonymous):

So -2 is just the answer? & Thank you.

OpenStudy (mertsj):

yes